If an electron and proton have the same de broglie wavelength which particle have greater speed

Grade:12

1 Answers

sonika p
6 years ago
de Broglie wavelength lambda = h/mv
for proton, lambda(p) = h/(mp)(vp) where, mp,vp is mass,velocity of proton,
for proton, lambda(e) = h/(me)(ve) where, me,ve is mass,velocity of electron, 
now when wavelengths are same,  velocity is inversely proportional to masses
Thus, ve>vp since, me
thus electron will have grater speed than proton.